# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v2

@@ 3.2.1 Controlling the Level of Fairness
3.3 Evaluation
As mentioned in Section 3, because we are dealing with stochastic retrievers, we need to measure the expected behavior of the system. Let $\mathcal{S}(\mathbf{s},N,k)$ be the stochastic sampler that samples a set of $N$ rankings $\sigma=\{\pi\}$, given the retrieval scores $\mathbf{s}$, where each ranking $\pi$ is truncated to the size of $k$. From each ranking, we can get an output $\hat{y}=\mathcal{G}(\phi_{p}(x,\pi))$. With an arbitrary fairness metric $\mu_{f}(\sigma)$ and a ranking quality metric $\mu_{r}(\sigma)$ that takes a set of rankings as an input, we can measure the degree of fairness and ranking quality of the sampled rankings. Similarly, an arbitrary string utility metric $\mu_{u}(y,\hat{y})$, such as ROUGE, can be used to assess an expected effectiveness of a RAG system by calculating the average of the $N$ metric scores.
In this paper, based on the empirical investigation done by Raj and Ekstrand [39], we use expected exposure disparity (EE-D) and expected exposure relevance (EE-R) [10] as $\mu_{f}$ and $\mu_{r}$, respectively (§3.3.1). For $\mu_{u}$, we select the metric depending on the task, and we get the expectation of the utility of a RAG model which we call an expected utility (EU) (§3.3.2).
3.3.1 Expected Exposure in the Context of Machine Users
Expected Exposure (EE) [10] works by estimating the exposure of items across rankings (e.g., $\sigma$) created by a subject model, and comparing them with an optimal set of rankings that always satisfy the item-fairness. To represent the attention over $n$ items given by the consumer (generator in RAG), an $n\times 1$ system exposure vector $\mathbf{\epsilon}$ is created. This is then compared with an $n\times 1$ target exposure vector $\mathbf{\epsilon}^{*}$, where it represents the exposure of items allocated by an oracle retriever that always rank useful items above non-useful ones [10].
@@ 3.3.1 Expected Exposure in the Context of Machine Users
${\left\lVert\mathbf{\epsilon}-\mathbf{\epsilon}^{*}\right\rVert}_{2}^{2}={\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}-2\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle+{\left\lVert\mathbf{\epsilon}^{*}\right\rVert}_{2}^{2}$ (3)
@@
The exposure of an item is calculated by modeling users’ (e.g., generators in RAG) attention to each item in a ranking. For example, one can assume that the user is affected by position bias and gives attention following an exponential decay [34]. However, these browsing models were developed for human-users not for machine-users, so we need a different user behavior model for generators in RAG. For simplicity, we assume that the machine-user can consume the items by giving equal attention to all the items that were passed to the context, but pays 0 attention to the items placed after the $k$’th position due to the top-$k$ truncation. This makes the user browsing model a step function parameterized by $k$. In this work, a relevance-independent machine-user model (MU) is set to the step function that reflects the behavior of top-k truncation of RAG:
$\text{MU}(i)=\begin{cases}1&\text{if }i\leq k\\ 0&\text{otherwise}\end{cases}=\mathds{1}[i\leq k]$ (4)
@@
$\mathbf{\epsilon}_{d}=\sum_{\pi\in S_{n}}p(\pi|q)\text{MU}(\bar{\pi}_{d})$ (5)
@@
$\mathbf{\epsilon}^{*}_{d}$ $=\frac{1}{m}\sum^{m}_{i=1}\text{MU}(i)=\begin{cases}1&\text{if }m\leq k\\ \frac{k}{m}&\text{otherwise}\end{cases}\quad\quad\mathbf{\epsilon}^{*}_{d^{-}}=\begin{cases}\frac{k-m}{n-m}&\text{if }m\leq k\\ 0&\text{otherwise}\end{cases}$ (6)
3.3.2 Expected Utility
@@
$\text{EU}(\langle\mathcal{S},\mathcal{G}\rangle|x)=\mathbb{E}_{\pi\sim\mathcal{S}}[\mu_{u}(y,\hat{y}_{\pi})]=\sum_{\pi\in S_{n}}p(\pi|q)\mu_{u}(y,\hat{y}_{\pi})\approx\frac{1}{N}\sum_{\pi\in\sigma}\mu_{u}(y,\hat{y}_{\pi})$ (7)
@@
3.3.3 Normalization of Metrics
From Equation 3, we decompose the metric into EE-D and EE-R. Since the bounds of these metrics depend on the number of useful items, normalization must be applied per query. Both metrics are min-max normalized based on their theoretical lower and upper bounds. We denote the normalized EE-D and EE-R as $\overline{\text{EE-D}}_{q}$ and $\overline{\text{EE-R}}_{q}$, respectively.
@@ 3.3.3 Normalization of Metrics
However, theoretically determining the bounds of the expected utility (EU) of a RAG model is challenging. To address this, we normalized the EU by the model’s empirical upper bound, the maximum observed utility across all runs of the experiment with the same generators. To approach the true upper bound, these runs include RAG models with an oracle retriever that consistently ranks useful items (i.e., those with positive utility labels) above non-useful ones, stochastically returning one of the $m!(n-m)!$ different rankings, where $m$ represents the number of useful items in the corpus. We denote the normalized EU as $\overline{\text{EU}}_{q}$, which can be interpreted as the distance to the optimal utility. From this section onwards, we omit the symbol $q$ from the normalized metrics for brevity. Proofs and details on how each metric is normalized by its lower and upper bounds can be found in the Appendix. C.
4 Experiment setup
@@
We choose the LaMP benchmark [46] for our dataset. It assesses the personalization capability of language models through retrieval-augmentation of users’ interaction history in a platform. LaMP includes various prediction tasks, such as classification, regression, and generation, and is well-suited for tasks where multiple items can be relevant/useful, unlike QA tasks with typically one or two provenance items. The retrieval items in LaMP have clear providers and consumers, aligning with our goal to ensure fairness for individual item providers. For example, in LaMP-1, retrieval items are academic papers, where exposure can increase citation counts for authors. In LaMP-4, retrieval items are news articles, where exposure can lead to monetary compensation for journalists. Due to the absence of a test set, we constructed a test collection as described in §3.1, using the first thousand entries of a user-based development set. Then, we discarded entries that have only one useful item in the corpus, as it is unnecessary to concern item-fairness in that case. We release the test collection, and the dataset statistics can be found in the Appendix J.
@@
Figure 3: Effect of a fairness control parameter $\alpha$ on the disparity of rankings, in the LaMP task 4, a text generation task. The RAG model is configured with the Contriever and Flan-T5-Base. Each data point represents the normalized EE-D of each run of the experiment (i.e., one query –>$N$ sampled rankings –> $\overline{\text{EE-D}}$ of the $N$ rankings).
@@
We use BM25 (lexical retriever) [40], SPLADE (learned sparse retriever), and Contriever (bi-encoder dense retriever) [21] as deterministic retrievers providing retrieval scores to base the sampling on. These models represent commonly used retrievers in the RAG literature [27]. We use a sampling size of $N=100$ and a truncation size of $k=5$.
@@
For generation models, we use Flan-T5-Small, Flan-T5-Base, and Flan-T5-XXL [8]. For decoding strategy, beam size is set to 4, and no sampling strategy is used. This is to ensure that stochasticity is only introduced to the retriever for controlled experiments. Full implementation details can be found in Appendix K. With the three base retrievers and three generators, we configure nine different RAG models and evaluate them on the seven LaMP tasks. Utility measurement of the generated strings followed the metrics used in the LaMP paper.
@@
We repeat the experiments with four different fairness control parameters $\alpha=1,2,4,8$, which allows us to assess the utility of the RAG models with different levels of item-fairness. From Figure 3, we observe how effectively $\alpha$, described in the Equation 2, controls the disparity of rankings. For example, when $\alpha$ is set to 4, we usually obtain a set of sampled rankings with $\overline{\text{EE-D}}$ mostly in the range of [0.5, 0.8], and when $\alpha$ is set to 8, we often get a set of sampled rankings with $\overline{\text{EE-D}}=1$. Refer to Appendix D to see the full description of the effect of $\alpha$ on the other metrics.
5 Results
Figure 4: Relationship between item-fairness and the quality of rankings.
@@ 5 Results
By gathering all four repeated runs of the experiments with different $\alpha$ values, we can plot the trend of ranking quality ($\overline{\text{EE-R}}$) against item fairness ($\overline{\text{EE-D}}$), as shown in Figure 4.
As shown in previous studies [51, 10], there is a well-known tradeoff between fairness and ranking quality for human users. Similarly, we observe a general tradeoff for machine users. However, unlike past findings, this tradeoff is not always strict. For instance, in Figure 4, both SPLADE and Contriever maintain consistently high ranking quality while being considerably fairer, and for BM25, ranking quality even improves as fairness increases, up to a certain point.
@@
The leftmost side of the lines, where $\overline{\text{EE-D}}=0$, represents the performance of a uniformly random ranking policy. At this point, the measured ranking quality should approximate the proportion of positively labeled items in the corpus, which is 31% based on data statistics (Appendix J). This is notably higher than in non-RAG (human-user) settings, where the percentage of relevant documents is typically much smaller, resulting in a $\overline{\text{EE-R}}$ value near 0 [10].
To quantify the performance of fair rankers, we calculate the area under the disparity-ranking quality curve (Figure 4), with higher values indicating stronger ranking quality. We also measure the tradeoff by fitting a linear line to the experiment results, where a steeper slope reflects a stronger tradeoff between fairness and ranking quality. Based on these metrics, we observe that Contriever-based models exhibit the highest tradeoff, while BM25-based models show the lowest, despite their poor retrieval quality. Overall, SPLADE-based models achieve high retrieval quality while maintaining a relatively low tradeoff. For detailed plots and quantifications, refer to Appendix F.
(a) Ranking Quality Vs. Utility
(b) Item-Fairness Vs. Utility
Figure 5: (a) Strong correlation between ranking quality and the generation quality of RAG. (b) Fairness-generation quality tradeoff. Full plots and quantifications are provided in Appendix G and H.
Fairness Intervals
@@
BM25+FlanT5Base (0.670) -0.20 -0.04 -0.08 -0.05 -0.02
BM25+FlanT5XXL (0.531) -0.07 +0.03 +0.02 +0.06 +0.11
@@
SPLADE+FlanT5Base (0.646) -0.15 +0.06 +0.08 0.00 +0.03
SPLADE+FlanT5XXL (0.671) -0.18 -0.16 +0.05 +0.02 +0.01
@@
Contriever+FlanT5Base (0.637) -0.16 +0.05 -0.06 +0.03 0.00
Contriever+FlanT5XXL (0.651) -0.19 -0.04 -0.11 +0.03 0.00
Fairness Intervals
@@
BM25+FlanT5Small (0.217) -0.06 0.00 +0.02 +0.01 0.00
BM25+FlanT5Base (0.223) -0.06 0.00 +0.03 +0.01 +0.02
BM25+FlanT5XXL (0.322) -0.05 +0.11 +0.03 +0.03 +0.05
SPLADE+FlanT5Small (0.235) -0.07 -0.01 +0.02 +0.03 +0.02
SPLADE+FlanT5Base (0.268) -0.10 -0.03 +0.02 0.00 +0.02
SPLADE+FlanT5XXL (0.342) -0.06 +0.09 +0.05 +0.03 +0.04
Contriever+FlanT5Small (0.254) -0.09 -0.02 0.00 +0.01 0.00
Contriever+FlanT5Base (0.268) -0.10 -0.02 +0.01 0.00 +0.01
Contriever+FlanT5XXL (0.367) -0.09 +0.06 +0.01 +0.01 +0.03
Table 1: Each value in the table is the difference between the utility of a baseline (deterministic) RAG model and the average utility of a fairer RAG model at a specific fairness interval. Nonnegative differences are highlighted. Full results are listed in Appendix I.
@@
Before examining the relationship between fairness and RAG utility, Figure 5(a) shows an auxiliary result confirming a strong correlation between utility-based ranking quality and the effectiveness of RAG models. This is unsurprising, as item-worthiness judgments were based on the utility-gain provided by the generator. However, this correlation suggests that the tradeoff observed in the disparity-ranking quality curve (Figure 4) is likely to manifest similarly due to this strong relationship.
@@
In fact, as observed from the disparity-utility curve (Figure 5(b)), we see a global trend of a non-strict tradeoff (i.e., RAG models maintain high generation quality while being considerably fair, and often even achieve higher quality).
@@
However, a closer look at the local trend offers a significant insight: RAG systems with fair ranking can often achieve higher system-effectiveness compared to models with deterministic rankers. In Table 1, we divided the fairness levels into five intervals based on the normalized EE-D. As shown in the table and Appendix I, improving fairness to the level of $\overline{\text{EE-D}}\in[0.8,1.0)$, and even $\overline{\text{EE-D}}\in[0.6,0.8)$, can often enhance the utility of many RAG models across most LaMP tasks. For example, having $\overline{\text{EE-D}}$ in the range of [0.8, 1.0) outperforms the baseline for all models in LaMP-2 and for seven out of nine models in LaMP tasks 4, 5, and 6.
6 Discussion and Conclusion

### v3

@@ 3.2.1. Controlling the Level of Fairness
3.3. Evaluation
As mentioned in Section 3, because we are dealing with stochastic retrievers, we need to measure the expected behavior of the system. Let $\mathcal{S}(\mathbf{s},N,k)$ be the stochastic sampler that samples a set of $N$ rankings $\sigma=\{\pi\}$, given the deterministic retrieval scores $\mathbf{s}$, where each ranking $\pi$ is truncated to the size of $k$. From each ranking, we can get an output $\hat{y}_{\pi}=\mathcal{G}(\phi_{p}(x,\pi))$. With an arbitrary fairness metric $\mu_{f}(\sigma)$ and a ranking quality metric $\mu_{r}(\sigma)$ that takes a set of rankings as an input, we can measure the degree of fairness and ranking quality of the sampled rankings. Similarly, an arbitrary string utility metric $\mu_{u}(y,\hat{y}_{\pi})$, such as ROUGE, can be used to assess an expected effectiveness of a RAG system by calculating the average of the $N$ metric scores.
In this paper, based on the empirical investigation done by Raj and Ekstrand (2020), we use expected exposure disparity (EE-D) and expected exposure relevance (EE-R) (Diaz et al., 2020) as $\mu_{f}$ and $\mu_{r}$, respectively (§3.3.1). For $\mu_{u}$, we select the metric depending on the task, and we get the expectation of the utility of a RAG model which we call an expected utility (EU) (§3.3.2). Beyond these metrics, we also measure the expected rate of attributed items (EAR) in §3.3.3, capturing how many retrieved items are ultimately used by the generator, and introduce the expected attributed exposure (EAE) in §3.3.4, which evaluates fairness specifically among the items that are actually attributed in the final output that can ultimately be displayed to human users.
3.3.1. Expected Exposure in the Context of Machine Users
Expected Exposure (EE) (Diaz et al., 2020) works by estimating the exposure of items across rankings (e.g., $\sigma$) created by a subject model, and comparing them with an optimal set of rankings that always satisfy the item-fairness. To represent the attention over $n$ items given by the consumer (generator in RAG), an $n\times 1$ system exposure vector $\mathbf{\epsilon}$ is created. This is then compared with an $n\times 1$ target exposure vector $\mathbf{\epsilon}^{*}$, where it represents the exposure of items allocated by an oracle retriever that always rank useful items above non-useful ones (Diaz et al., 2020).
@@ 3.3.1. Expected Exposure in the Context of Machine Users
(3) ${\left\lVert\mathbf{\epsilon}-\mathbf{\epsilon}^{*}\right\rVert}_{2}^{2}={\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}-2\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle+{\left\lVert\mathbf{\epsilon}^{*}\right\rVert}_{2}^{2}$
@@
The exposure of an item is calculated by modeling users’ (e.g., generators in RAG) attention to each item in a ranking. For example, one can assume that the user is affected by position bias and gives attention following an exponential decay (Moffat and Zobel, 2008). However, these browsing models were developed for human-users not for machine-users, so we need a different user behavior model for generators in RAG. For the simplicity of the metric, and given recent efforts to reduce position bias and promote more even attention in machine-user settings (He et al., 2024; Hsieh et al., 2024), we assume that the machine user consumes all items passed to the context with equal attention but pays zero attention to items placed after the $k$’th position due to top-$k$ truncation. This makes the user browsing model a step function parameterized by $k$. In this work, a relevance-independent machine-user model (MU) is set to the step function that reflects the behavior of top-k truncation of RAG:
(4) $\text{MU}(i)=\begin{cases}1&\text{if }i\leq k\\ 0&\text{otherwise}\end{cases}=\mathds{1}[i\leq k]$
@@
(5) $\mathbf{\epsilon}_{d}=\sum_{\pi\in S_{n}}p(\pi|q)\text{MU}(\bar{\pi}_{d})$
@@
(6) $\mathbf{\epsilon}^{*}_{d}$ $=\frac{1}{m}\sum^{m}_{i=1}\text{MU}(i)=\begin{cases}1&\text{if }m\leq k\\ \frac{k}{m}&\text{otherwise}\end{cases}\quad\mathbf{\epsilon}^{*}_{d^{-}}=\begin{cases}\frac{k-m}{n-m}&\text{if }m\leq k\\ 0&\text{otherwise}\end{cases}$
3.3.2. Expected Utility
@@
(7) $\text{EU}(\langle\mathcal{S},\mathcal{G}\rangle|x)$ $=\mathbb{E}_{\pi\sim\mathcal{S}}[\mu_{u}(y,\hat{y}_{\pi})]$
$=\sum_{\pi\in S_{n}}p(\pi|q)\mu_{u}(y,\hat{y}_{\pi})\approx\frac{1}{N}\sum_{\pi\in\sigma}\mu_{u}(y,\hat{y}_{\pi})$
@@
3.3.3. Expected Attribution Rate
We use a natural language inference (NLI) model for attribution, reflecting its effectiveness in measuring faithfulness. Honovich et al. (2022) find that NLI outperforms other metrics in a meta-evaluation, leading works such as automated citation evaluation using NLI (Gao et al., 2023b), which strongly correlates with human judgments. Building on the human evaluation framework by Rashkin et al. (2023), Gao et al. (2023a) introduced a new NLI-based metric to approximate human judgments, followed by Bohnet et al. (2022). These studies collectively show that NLI-based methods capture attribution quality in a manner closely aligned with human assessments, making them a reliable choice for our evaluation.
Given a ranking $\pi\in\sigma$ of $k$ items and a predicted output $\hat{y}_{\pi}$, we therefore measure item attribution rate (AR) using
(8) $\mu_{\text{AR}}\bigl{(}\pi,\hat{y}_{\pi}\bigr{)}\;=\;\frac{1}{k}\sum_{d\in\pi}\text{NLI}(d,\hat{y}_{\pi})$
where $\text{NLI}(d,\hat{y}_{\pi})\!=\!1$ if item $d$ entails the output $\hat{y}_{\pi}$, and 00 otherwise, by a natural language inference model. 444 This metric is similar to the context utilization metric from Ru et al. (2024), which also measures how effectively the generator leverages retrieved context. However, Ru et al. (2024) decompose the generated response into multiple claims and focus only on the proportion of relevant items, whereas our approach applies a single NLI check for each retrieved item.
Analogous to expected utility, we define expected attribution rate (EAR) of a RAG system as
(9) $\text{EAR}(\langle\mathcal{S},\mathcal{G}\rangle|x)$ $=\mathbb{E}_{\pi\sim\mathcal{S}}[\mu_{\text{AR}}(\pi,\hat{y}_{\pi})]$
$=\sum_{\pi\in S_{n}}p(\pi|q)\mu_{\text{AR}}(\pi,\hat{y}_{\pi})\approx\frac{1}{N}\sum_{\pi\in\sigma}\mu_{\text{AR}}(\pi,\hat{y}_{\pi}).$
3.3.4. Expected Attributed Exposure
While EE measures retrieval fairness and relevance, it does not capture whether the retrieved items actually appear in the final generated output or how fairly exposure is allocated among those items once they surface after generation. To address this gap, we define the attributed exposure vector $\mathbf{\epsilon}^{a}\in\mathbb{R}^{n}$, where an attributed exposure for each item $d$ is calculated as
(10) $\mathbf{\epsilon}^{a}_{d}=\sum_{\pi\in S_{n}}p(\pi|q)\text{NLI}(d,\hat{y}_{\pi}).$
Similar to EE, a disparity measure expected attributed exposure disparity (EAE-D) can be derived by measuring the squared $l_{2}$ distance of an attributed exposure vector, thus
[…]
Hence, EAE-D measures how fairly exposure is allocated across the items that are explicitly used, addressing the shortcoming of EE by directly linking exposure to items that influence the final output of a RAG system.
3.3.5. Normalization of Metrics
Since EAR is an average over the percentage of entailed items for each sampled ranking, its values lie on a consistent scale with [0, 1] range. However, the bounds of EE-D, EE-R, and EAE-D depend on the number of useful items in the corpus. Consequently, we apply normalization on a per-query basis by min-max scaling each metric according to its theoretical lower and upper bounds. We denote the normalized EE-D, EE-R, and EAE-D as $\overline{\text{EE-D}}$, $\overline{\text{EE-R}}$, and $\overline{\text{EAE-D}}$ respectively.
@@ 3.3.5. Normalization of Metrics
However, theoretically determining the bounds of the expected utility (EU) of a RAG model is challenging. To address this, we normalized the EU by the model’s empirical upper bound, the maximum observed utility across all runs of the experiment with the same generators. To approach the true upper bound, these runs include RAG models with an oracle retriever that consistently ranks useful items (i.e., those with positive utility labels) above non-useful ones, stochastically returning one of the $m!(n-m)!$ different rankings, where $m$ represents the number of useful items in the corpus. We denote the normalized EU as $\overline{\text{EU}}$, which can be interpreted as the distance to the optimal utility.
Proofs and details on how each metric is normalized by its lower and upper bounds can be found in the Appendix C.
4. Experiments
@@
We choose the LaMP benchmark (Salemi et al., 2024b) for our dataset. It assesses the personalization capability of language models through retrieval-augmentation of users’ interaction history in a platform. LaMP includes various prediction tasks, such as classification, regression, and generation, and is well-suited for tasks where multiple items can be relevant/useful, unlike QA tasks with typically one or two provenance items. The retrieval items in LaMP have clear providers and consumers, aligning with our goal to ensure fairness for individual item providers. For example, in LaMP-1, retrieval items are academic papers, where exposure can increase citation counts for authors. In LaMP-4, retrieval items are news articles, where exposure can lead to monetary compensation for journalists.  [^5] Throughout this paper, we use LaMP-1 to represent classification and LaMP-4 to represent text generation in visualizations. All LaMP tasks were used in our experiments, and similar trends were observed. Due to the absence of a test set, we constructed a test collection as described in §3.1, using the first thousand entries of a user-based development set. Then, we discarded entries that have only one useful item in the corpus, as it is unnecessary to concern item-fairness in that case. We release the test collection along with dataset statistics. Dataset statistics can be found in the Appendix E.
(a) LaMP 1
(b) LaMP 4
@@
Figure 3. Effect of a fairness control parameter ($\alpha$) on the disparity of rankings ($\overline{\text{EE-D}}$) in LaMP task 1 and 4. The RAG model is configured with the Contriever and Flan-UL2. Each data point represents the normalized EE-D of each run of the experiment (i.e., one query –¿$N$ sampled rankings –¿ $\overline{\text{EE-D}}$ of the $N$ rankings).
@@
We use BM25 (lexical retriever) (Robertson et al., 1995), SPLADE (learned sparse retriever) (Formal et al., 2022), and Contriever (bi-encoder dense retriever) (Izacard et al., 2022) as deterministic retrievers providing retrieval scores to base the sampling on. These models represent commonly used retrievers in the RAG literature (Kim et al., 2024). We use a sampling size of $N=100$ and a truncation size of $k=5$.
@@
For generation models, we use Flan-T5-Small, Flan-T5-Base, Flan-T5-XXL (Chung et al., 2024), and Flan-UL2 (Tay et al., 2023). For decoding strategy, beam size is set to 4, and no sampling strategy is used. This is to ensure that stochasticity is only introduced to the retriever for controlled experiments. With the three base retrievers and four generators, we configure twelve different RAG models and evaluate them on the seven LaMP tasks. Utility measurement of the generated strings follows the metrics used in the LaMP paper. For NLI computations, a RoBERTa large model (Liu, 2019) fine-tuned on the natural language inference task is used. [^6] https://huggingface.co/FacebookAI/roberta-large-mnli
@@
We repeat the experiments with four different fairness control parameters $\alpha=1,2,4,8$, which allows us to assess the utility of the RAG models with different levels of item-fairness. From Figure 3, we observe how effectively $\alpha$, described in the Equation 2, controls the disparity of rankings. For example, when $\alpha$ is set to 4, we usually obtain a set of sampled rankings with $\overline{\text{EE-D}}$ mostly in the range of [0.5, 0.8], and when $\alpha$ is set to 8, we often get a set of sampled rankings with $\overline{\text{EE-D}}=1$.
We conducted paired (per-query) $t$-tests for $\overline{\text{EE-D}}$ across the different $\alpha$ values and found statistically significant differences $(p<0.01)$ in all 84 experimental conditions ($12$ models $\times$ $7$ tasks).
5. Results
@@
By gathering all four repeated runs of the experiments with different $\alpha$ values, we can plot the trend of ranking quality ($\overline{\text{EE-R}}$) against item fairness ($\overline{\text{EE-D}}$), as shown in Figure 4(a).
As shown in previous studies (Singh and Joachims, 2019; Diaz et al., 2020), there is a well-known tradeoff between fairness and ranking quality for human users. Similarly, we observe a general tradeoff for machine users. However, unlike past findings, this tradeoff is not always strict. For instance, in Figure 4(a), both SPLADE and Contriever maintain consistently high ranking quality while being considerably fairer, and for BM25, ranking quality even improves as fairness increases, up to a certain point.
@@
The leftmost side of the lines, where $\overline{\text{EE-D}}=0$, represents the performance of a uniformly random ranking policy. At this point, the measured ranking quality should approximate the proportion of positively labeled items in the corpus, which is 31% for LaMP task 4. This is notably higher than in non-RAG (human-user) settings, where the percentage of relevant documents is typically much smaller, resulting in a $\overline{\text{EE-R}}$ value near 0 (Diaz et al., 2020).
[…]
BM25 0.1351 0.3603
SPLADE 0.1971 0.4166
Contriever 0.1741 0.3864
(a) LaMP-1
[…]
BM25 0.0895 0.3921
SPLADE 0.1453 0.4205
Contriever 0.1624 0.4287
(b) LaMP-4
Table 1. Comparison of fairness-ranking quality tradeoff (slope) and ranking quality (AUC) metrics for LaMP-1 and LaMP-4 datasets.
To quantify the tradeoff, we fit a linear line to the experiment results, where a steeper slope reflects a stronger tradeoff between fairness and ranking quality (slope). We also quantify the performance of fair rankers, by calculating the area under the disparity-ranking quality curve (AUC; Figure 4(a)). As can be seen in Table 1, we observe that retriever that yields higher ranking quality has higher tradeoff in terms of retrieval fairness.
(a) RetDisparity vs. Ranking Quality
(b) Ranking Quality vs. Utility
(c) RetDisparity vs. Utility
Figure 4. Relationships between retrieval fairness, ranking quality, and generation quality. We empirically show that there is a positive relationship between ranking quality and utility, but trade-offs between retrieval fairness and ranking quality, as well as retrieval fairness and utility in RAG models. The curves are plotted by interpolating the runs for each disparity interval and averaging across generators for 4(a) and across retrievers for 4(b) and 4(c). RetDisparity refers to the disparity among retrieved items.
Disparity Intervals
@@
BM25+FlanT5Base (0.670) -0.20* -0.04 -0.08 -0.05 -0.02
BM25+FlanT5XXL (0.531) -0.07* +0.03 +0.02 +0.06 +0.11*
BM25+FlanUL2 (0.557) -0.1* -0.02 -0.02 +0.01 +0.03
@@
SPLADE+FlanT5Base (0.646) -0.15* +0.06 +0.08 0.00 +0.03
SPLADE+FlanT5XXL (0.671) -0.18* -0.16 +0.05 +0.02 +0.01
SPLADE+FlanUL2 (0.632) -0.15* -0.08 +0.02 +0.02 +0.01
@@
Contriever+FlanT5Base (0.637) -0.16* +0.05 -0.06 +0.03 0.00
Contriever+FlanT5XXL (0.651) -0.19* -0.04 -0.11 +0.03 0.00
Contriever+FlanUL2 (0.620) -0.16* -0.09 -0.05 +0.02 -0.03
Disparity Intervals
@@
BM25+FlanT5Small (0.217) -0.06* 0.00 +0.02 +0.01 0.00
BM25+FlanT5Base (0.223) -0.06* 0.00 +0.03 +0.01 +0.02
BM25+FlanT5XXL (0.322) -0.05* +0.11* +0.03 +0.03 +0.05*
BM25+FlanUL2 (0.381) -0.07* +0.06* +0.07* +0.01 +0.04
SPLADE+FlanT5Small (0.235) -0.07* -0.01 +0.02 +0.03* +0.02
SPLADE+FlanT5Base (0.268) -0.10* -0.03 +0.02 0.00 +0.02
SPLADE+FlanT5XXL (0.342) -0.06* +0.09* +0.05* +0.03 +0.04*
SPLADE+FlanUL2 (0.429) -0.12* +0.04 +0.02 +0.01 +0.01
Contriever+FlanT5Small (0.254) -0.09* -0.02* 0.00 +0.01 0.00
Contriever+FlanT5Base (0.268) -0.10* -0.02 +0.01 0.00 +0.01
Contriever+FlanT5XXL (0.367) -0.09* +0.06* +0.01 +0.01 +0.03
Contriever+FlanUL2 (0.449) -0.15* +0.01 -0.01 0.00 +0.02
Table 2. Each value in the table is the difference between the utility of a baseline (deterministic) RAG model and the average utility of a fairer RAG model at a specific retrieval disparity interval. Nonnegative differences are highlighted. All LaMP tasks shows similar trend. A superscript * denotes a statistically significant difference between the treatment performance and the baseline performance based on an unpaired Student’s t-test ($p<0.05$). Full results are listed in Appendix D.
@@
Before examining the relationship between fairness and RAG utility, Figure 4(b) shows an auxiliary result confirming a strong correlation between utility-based ranking quality and the effectiveness of RAG models. This is unsurprising, as item-worthiness judgments were based on the utility-gain provided by the generator. However, this correlation suggests that the tradeoff observed in the disparity-ranking quality curve (Figure 4(a)) is likely to manifest similarly due to this strong relationship.
@@
In fact, as observed from the disparity-utility curve (Figure 4(c)), we see a global trend of a non-strict tradeoff (i.e., RAG models maintain high generation quality while being considerably fair, and often even achieve higher quality).
@@
However, a closer look at the local trend offers a significant insight: RAG systems with fair ranking can often achieve higher system-effectiveness compared to models with deterministic rankers. In Table 2, we divided the fairness levels into five disparity intervals based on the normalized EE-D. As shown in the table , improving fairness to the level of $\overline{\text{EE-D}}\in[0.8,1.0)$, and even $\overline{\text{EE-D}}\in[0.6,0.8)$, can often enhance the expected utility of many RAG models across LaMP tasks. For example, having $\overline{\text{EE-D}}$ in the range of [0.8, 1.0) outperforms the baseline for all models in LaMP-4 and for eight out of twelve models in LaMP-1.
Significance testing results in Table 2 provide additional nuance. The range $\overline{\text{EE-D}}\in[0.0,0.2)$ shows a statistically significant drop in utility compared to the baseline, indicating that pushing disparity to extremely low levels can indeed reduce performance. Beyond this small-disparity range ($\overline{\text{EE-D}}\in[0.2,1.0)$), the differences in utility scores from the baseline either remain statistically insignificant or—when significant—reflect improved utility. This suggests that models can maintain a level of retrieval fairness close to the baseline’s utility performance without incurring a notable cost, and may even achieve higher effectiveness in many cases.
(a) RetDisparity vs. Attribution Rate
(b) Ranking Quality vs. Attribution Rate
(c) RetDisparity vs. AttDisparity
Figure 5. Major relationships between retrieval fairness (RetDisparity), ranking quality, expected attribution rate, and consumption fairness (AttDisparity). We empirically show that there is a tradeoff between retrieval fairness and expected attribution rate as there is a positive relationship between ranking quality and attribution rate. Most importantly, we show that retrieval fairness does not necessarily directly propagates to the consumption fairness due to varying attribution rate of a generator. The same interpolation methods are used as Figure 4 and runs were averaged across generators. RetDisparity refers to the disparity among retrieved items, while AttDisparity refers to the disparity among attributed sources.
RQ3: What is the impact of item-fairness in ranking to the fairness of the attributed sources used in the final response?
To address this, we analyze three primary relationships: (1) how retrieval fairness ($\overline{\text{EE-D}}$) relates to the expected attribution rate (EAR), (2) how ranking quality ($\overline{\text{EE-R}}$) interacts with EAR, and (3) how retrieval fairness translates into consumption fairness ($\overline{\text{EAE-D}}$) in the final output.
Figures 5(a) and 5(b) together shed light on how many items are actually attributed by the generator, which is a vital step in understanding how exposure might become more or less equitable once the generator has consumed the retrieved items. Figure 5(a) shows that when retrieval becomes more fair (lower $\overline{\text{EE-D}}$), the generator tends to attribute fewer items overall, suggesting a tradeoff between distributing items equitably and actually getting them used in the final output.
Interestingly, the general shape of $\overline{\text{EE-D}}$ versus EAR in Figure 5(a) aligns closely with the shape of $\overline{\text{EE-D}}$ versus $\overline{\text{EE-R}}$ in Figure 4(a). By looking at Figure 5(b), it becomes clear that the generator is more likely to use multiple items if they are useful in solving the task, reflecting the idea that high-quality rankings encourage higher EAR.
Figure 5(c) directly addresses the question of whether fair retrieval leads to fair consumption by comparing $\overline{\text{EE-D}}$ (retrieval fairness) to $\overline{\text{EAE-D}}$ (consumption fairness).  [^7] Normalized EE-D and EAE-D scores are directly comparable, as they share the same dimensions in their exposure vectors and have the same $l1$ norm. The positive trend suggests that a more equitable ranking generally produces a more equitable distribution of attributed items in the final generation. However, the diagonal line (y=x) reveals subtle differences between fairness at the ranking stage and fairness at the consumption stage. Data points falling below this line imply that consumption is even fairer than the initial ranking might predict, while points above it suggest that the generator’s selective usage of items has introduced new disparity. This pattern shifts depending on how relevant and useful the retrieved items are, mirroring the earlier observation (Figure 5(b)) that the generator favors content that supports solving the task. In cases where many items have only marginal utility, the generator’s emphasis on a few high-utility sources can inadvertently raise disparity, placing the plotted values above the y=x line.
6. Discussion
6.1. Higher System Utility with Fair Rankings